# ComKD-CLIP: Comprehensive Knowledge Distillation for Contrastive Language-Image Pre-traning Model

### v1

@@ Conclusion
In this study, we present ComKD-CLIP: Comprehensive Knowledge Distillation for the Contrastive Language-Image Pre-training (CLIP) model. This innovative method employs ImgAlign and EduAttention to effectively distill the complex knowledge embedded in the large CLIP model during the text-image feature fusion process. Furthermore, ComKD-CLIP refines the distilled knowledge by harnessing the feature fusion results of the large model. This strategy enables the smaller model to assimilate the nuanced knowledge from the larger model, thereby equipping the smaller model with capabilities to achieve comparable performance. Extensive experiments across eleven datasets have unequivocally demonstrated the superior performance of the proposed method. ComKD-CLIP significantly enhances the performance of smaller CLIP models in resource-constrained environments and augments the practical applicability of CLIP technology.

### v3

@@ Conclusion
In this study, we present ComKD-CLIP, an advanced knowledge distillation framework designed to comprehensively distill the knowledge from a large teacher CLIP model to a smaller student model. This process maintains the comparable performance while substantially reducing the model’s parameter. ComKD-CLIP innovatively employs IFAlign and EduAttention to effectively distill the intricate knowledge embedded within the teacher CLIP model during the text-image feature fusion process. Futhermore, ComKD-CLIP can refine the distilled knowledge by leveraging the feature fusion results of the teacher model, ensuring that the smaller student model can accurately absorb the knowledge from teacher model. Extensive experiments across 11 datasets have unequivocally demonstrated the superior performance of ComKD-CLIP. The proposed method significantly bolsters the capabilities of smaller CLIP models in resource-constrained environments, thereby broadening the practical utility of the CLIP technology.